# The purpose of an estimator is what it does: Misspecification, estimands, and over-identification

### v4

@@ 5.1. J-Statistics, not J-Tests
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}\operatorname{P}\left(\left|\inf_{\theta_{0}\in\mathbb{R}}\sup_{\Omega\in\mathcal{W}_{\kappa}}|t_{\Omega}(\theta_{0})|-\sqrt{J}\right|>\varepsilon\right)=0.$
This result shows that, asymptotically, a researcher can find a weighting matrix that yields an absolute t-statistic of at least $\sqrt{J}$ for any possible null hypothesis. Hence, for instance, when $\sqrt{J}>1.96$ a nefarious researcher who is unconstrained in their choice of weighting matrix can reject any null hypothesis at the 5% level. Consequently, in order to limit the scope for “weight-hacking” to yield statistically significant results, it must be that either $J$ statistics are small or that other restrictions constrain the choice of weight matrix. [^26] Among other results, Honoré et al. (2020) characterize the derivative of $\sigma_{\Omega}^{2}$ with respect to elements of $\Omega.$ Their results may thus be useful for building intuition about which $\Omega$ satisfy $\hat{\sigma}_{\Omega}^{2}\leq(1+\tau^{2})\hat{\sigma}_{\Sigma^{-1}}^{2}$.
Absent restrictions on the weighting matrix the J-statistic threshold necessary for a result to be “weight-hacking” proof at the 5% level, $\sqrt{J}\leq 1.96$ does not scale with the degree of over-identification, and so can be much more stringent than conventional J-test critical values. In Table 1, for instance, four of ten specifications, and 3 of 6 papers, have $\sqrt{J}>1.96$, despite only one specification being rejected at the 5% level by the $J$-test.
@@
Under the assumptions of 5.2, for a critical value $c$, consider confidence sets centered around the GMM estimator $\hat{\theta}_{\Omega}$, $\hat{C}_{\Omega}(c)=[\hat{\theta}_{\Omega}\pm c\hat{\sigma}_{\Omega}]$. Then for all $\varepsilon>0,$
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}\operatorname{P}\left(\left|\inf\left\{{c\geq 0:\bigcap_{\Omega\in\mathcal{W}_{\kappa}}C_{\Omega}(c)\neq\varnothing}\right\}-\sqrt{J}\right|>\varepsilon\right)=0.$
@@
$\lim_{\nu\to 0}\lim_{\kappa\to\infty}\limsup_{n\to\infty}\operatorname{P}\left(\left|\sqrt{n}\cdot d_{H}\left(\bigcap_{\Omega\in\mathcal{W}_{\kappa}}C_{\Omega}(\sqrt{J}+\nu),\left\{\hat{\theta}_{\Sigma^{-1}}\right\}\right)\right|>\varepsilon\right)=0.$
Hence, to ensure at least one point of overlap between the confidence sets of researchers considering different weighting matrices, researchers would need to use critical values no smaller than $\sqrt{J}.$ This result bears an interesting resemblance to one by Andrews et al. (2024). There, working in a Bayesian framework and studying linear-in-parameters moment condition models, the authors show that confidence intervals with width $\sqrt{J}$ have correct average coverage under a class of priors motivated by the GMM criterion.
Our asymptotic results in this section follow from exact, finite-sample results in a model, discussed in the Appendix, where the moments are linear in parameters and have known variance. Armstrong and Kolesár (2021) show that this linear model corresponds to the limit experiment for weakly identified GMM.
Returning to Table 1, of the 10 specifications, 4 have J-statistics larger than $1.96^{2}=3.84$; for these specifications, it is possible to choose weight matrices to reject any point null hypothesis. More quantitatively, using 5.2, we can interpret conventional confidence sets around the efficient estimator $\hat{\theta}_{\Sigma^{-1}}$ as a set of GMM estimates at a given standard error inflation $\sqrt{1+\tau^{2}}-1$ for $\tau=1.96/\sqrt{J}$. The interquartile range of this implied standard error inflation is 35% to 230%.
@@
Other recent work includes Kitamura et al. (2013), who establish a local-misspecification-robustness property for a generalized empirical likelihood estimator, and Andrews et al. (2020) who use a local misspecification framework to examine the relationship between structural estimates and other, non-moment statistics, again with the goal of helping researchers clarify what “drives” their estimates. Fessler and Kasy (2019) show how one can use the implications of locally misspecified economic models to improve unbiased but potentially noisy estimates, while Armstrong et al. (2024) consider the problem of optimally combining biased and unbiased estimators under local misspecification. $\blacksquare$
@@ Appendix A Normal Limit Problem
Armstrong and Kolesár (2021) show that the problem of estimating $\theta$ under local misspecification reduces, asymptotically, to that of observing $(Y,\Gamma,\Sigma,h)$ where we wish to estimate $\theta=h^{\prime}\phi$ and
@@
Quantities associated with GMM have natural analogues in the limit problem (A), which we collect in Table 2. These quantities are linked in the sense that the asymptotic behavior of GMM quantities (left-side of Table 2) agree with the behavior of the limit problem analogues under (A). In particular, GMM estimators for $\psi$ with weight matrix $\Omega$ correspond to estimators of $\phi$ in the form
@@
$=(Y+\Gamma\hat{\phi}_{\Sigma^{-1}})\Sigma^{-1}(Y+\Gamma\hat{\phi}_{\Sigma^{-1}})$
@@
Jacobian of moments, $\mathbb{E}_{P}\left[{\frac{\partial{}}{\partial{\psi}}g(X_{i},\psi)}\right]$ $\Gamma$
@@
J-statistic $\min_{\phi\in\Psi}g_{n}(\phi)^{\prime}\operatorname{Var}_{P}(g(X_{i},\psi))^{-1}g_{n}(\phi)$ $\min_{u}(Y+\Gamma u)\Sigma^{-1}(Y+\Gamma u)$
@@
To further motivate the limit problem (A), we briefly sketch the connection to the asymptotics of GMM. We first localize the problem to a $1/\sqrt{n}$-neighborhood of $\psi$. Suppose we are given some (nonrandom) $\tilde{\psi}$ with
$\phi\equiv\sqrt{n}(\tilde{\psi}-\psi)=O(1).$
We can view $\phi$ as the parameter of interest instead. Any estimator $\hat{\psi}$ of $\psi$ induces an estimator $\hat{\phi}$ of $\phi$ since
@@
$t_{\Omega}(\theta_{0})=\frac{h^{\prime}\hat{\phi}_{\Omega}-\theta_{0}}{\sqrt{\operatorname{Var}(h^{\prime}\phi_{\Omega})}}.$
@@
Moreover, recall that $\hat{\theta}_{\Omega}=-h^{\prime}(\Gamma_{P}^{\prime}\Omega_{P}\Gamma)^{-1}\Gamma_{P}^{\prime}\Omega_{P}Y$, $\sigma_{\Omega}^{2}=\operatorname{Var}(\hat{\theta}_{\Omega})$, $\hat{\theta}_{\Sigma^{-1}}=\hat{\theta}_{\Sigma^{-1}}$, $\sigma_{\Sigma^{-1}}^{2}=\sigma_{\Sigma^{-1}}^{2}$. Then,
@@
    $\sigma^{2}_{\Omega}=\sigma_{\Sigma^{-1}}^{2}+v_{\Omega}^{\prime}\tilde{\Sigma}_{Z}^{*}v_{\Omega}$
@@
    The J-statistic is the norm of $Z$: $J=Z^{\prime}\left({\Sigma_{Z}^{*}}\right)^{-1}Z$.
@@ A.1. Proofs for Normal Limit Problem
As a result, for $v_{\Omega}=c{\Sigma^{*}_{Z}}^{-1}Z$, we have that $\hat{\theta}_{\Omega}=\hat{\theta}_{\Sigma^{-1}}+cJ$ and
$\sigma_{\Omega}^{2}=\sigma_{\Sigma^{-1}}^{2}+c^{2}J.$
@@
$\hat{\theta}_{\Omega}-\hat{\theta}_{\Sigma^{-1}}=\sigma_{\Sigma^{-1}}\tau\sqrt{J}\quad\sigma_{\Omega}=\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}$
@@
This is immediate from Corollary A.4. In particular, note that by Corollary A.4, $t_{\Omega}(\hat{\theta}_{\Sigma^{-1}})\leq\sqrt{J}$ for all $\Omega,$ while for any $\theta$ there exists $\Omega$ with $t_{\Omega}(\theta)\geq\sqrt{J}$. ∎
@@
We first show that
$\left\{{\Omega\Gamma(\Gamma^{\prime}\Omega\Gamma)^{-1}h:\Omega}\right\}$ $=\left\{{q\in\mathbb{R}^{k}:\Gamma^{\prime}q=h}\right\}$ (8)
$=\left\{{q\in\mathbb{R}^{k}:q=q_{0}+v,\Gamma^{\prime}q_{0}=h,v\in N(\Gamma^{\prime})}\right\}$
$q=u+w$
for $P_{V}$ the projection matrix onto $V$. Note that by construction, any vector $t$ can be written as
$t=\Gamma t_{1}+cw+t_{2}\quad t_{1}\in\mathbb{R}^{p}\quad t_{2}\in V$
Choose $a=\frac{1}{h^{\prime}(\Gamma^{\prime}\Gamma)^{-2}h}>0$. Observe that $\Omega$ is positive definite:
$\geq a(t_{1}^{\prime}h)^{2}+ac^{2}\lVert w\rVert^{4}+2ac\lVert w\rVert^{2}t_{1}^{\prime}h+\lVert t_{2}\rVert^{2}$ (By Cauchy–Schwarz, $(t_{1}^{\prime}h)^{2}\leq(t_{1}^{\prime}(\Gamma^{\prime}\Gamma)^{2}t_{1})(h^{\prime}(\Gamma^{\prime}\Gamma)^{-2}h)$ )
$=a(t_{1}^{\prime}h+c\lVert w\rVert^{2})^{2}+\lVert t_{2}\rVert^{2}\geq 0.$
Now,
and thus
This proves (8).
Thus, the set of $v_{\Omega}$, as $\Omega$ ranges over all weighting matrices, is the set
$\left\{{-(M\Sigma M^{\prime})^{-1}M\Sigma q_{0}-(M\Sigma M^{\prime})^{-1}M\Sigma v:v\in N(\Gamma^{\prime})}\right\}$
@@
$N(M\Sigma^{1/2})$ $=\left\{{v:M\Sigma^{1/2}v=0}\right\}=\left\{{v:\Sigma^{1/2}v\in R(\Gamma)}\right\}$ ($N(M)=R(\Gamma)$)
@@
$QY+Q\Gamma\hat{\phi}^{*}$ $=\begin{bmatrix}\hat{\phi}_{\Sigma^{-1}}\\ Z\end{bmatrix}-\begin{bmatrix}\hat{\phi}_{\Sigma^{-1}}\\ 0\end{bmatrix}=\begin{bmatrix}0\\ Z\end{bmatrix}$
@@ Appendix B Proofs for Results in Main Text
$\sqrt{n}(\hat{\theta}_{\Omega}-\vartheta(\psi))-\hat{\theta}^{*}_{\Omega}\overset{p}{\longrightarrow}0,~~\sqrt{n}\hat{\sigma}_{\Omega}\overset{p}{\longrightarrow}\sigma_{\Omega}=\left({h^{\prime}\left(\Gamma^{\prime}\Omega\Gamma\right)^{-1}\Gamma^{\prime}\Omega\Sigma\Omega\Gamma\left(\Gamma^{\prime}\Omega\Gamma\right)^{-1}h}\right)^{1/2},$ (9)
@@
$J=n\cdot g_{n}\left(\hat{\psi}_{\Sigma^{-1}}\right)^{\prime}\hat{\Sigma}g_{n}\left(\hat{\psi}_{\Sigma^{-1}}\right)$
@@
$J^{*}=\min_{u}(\sqrt{n}g_{n}(\psi)+\Gamma u)\Sigma^{-1}(\sqrt{n}g_{n}(\psi)+\Gamma u),$
@@
$\sqrt{n}\cdot\left(\hat{\Theta}(\tau)-\vartheta(\psi)\right)=\left[\sqrt{n}(\hat{\theta}_{\Sigma^{-1}}-\vartheta(\psi))\pm\tau\sqrt{J}\sqrt{n}\sigma_{\Sigma^{-1}}\right],$
@@
$\text{plim}\left(d_{H}\left(\sqrt{n}\cdot\left(\hat{\Theta}\left(\kappa,\tau\right)-\vartheta(\psi)\right),\hat{\Theta}^{*}\left(\kappa,\tau\right)\right)\right)=0.$ (10)
@@
then $\lim_{\delta\to 0}d_{H}\left(\mathcal{W}_{\kappa,\tau},\mathcal{W}_{\kappa,\tau+\delta}\right)=0$ . Consequently, since (9) implies that for $\delta>0$
$\operatorname{P}\left\{\mathcal{W}_{\kappa,\tau-\delta}\subseteq\mathcal{\hat{W}}_{\kappa,\tau}\subseteq\mathcal{W}_{\kappa,\tau+\delta}\right\}\to 1,$
@@
Our argument thus far implies that for each $\kappa>0$, and each $\nu>0,$
$\liminf_{n\to\infty}\operatorname{P}\left(\sqrt{n}\cdot d_{H}\left(\hat{\Theta}\left(\kappa,\tau\right),\hat{\Theta}(\tau)\right)\leq d_{H}\left(\hat{\Theta}^{*}\left(\kappa,\tau\right),\hat{\Theta}^{*}\left(\tau\right)\right)+\nu\right)=1.$
To complete the proof, we wish to show that the right hand side converges to zero as $\kappa\to\infty$. To this end, note that $\hat{\Theta}^{*}\left(\kappa,\tau\right)$ and $\hat{\Theta}^{*}\left(\tau\right)$ both depend on the data only through $\sqrt{n}g_{n}(\psi)$, and let
@@
$\liminf_{n\to\infty}\operatorname{P}\left\{d_{H}\left(\hat{\Theta}^{*}\left(\kappa,\tau\right),\hat{\Theta}^{*}\left(\tau\right)\right)<\zeta\right\}\geq\operatorname{P}\left\{\mathcal{N}\left(\eta,\Sigma\right)\in\mathcal{G}(\kappa,\zeta)\right\},$
@@
$\lim_{\kappa\to\infty}\operatorname{P}\left\{N\left(\eta,\Sigma\right)\in\mathcal{G}(\kappa,\zeta)\right\}=1,$
@@
$\inf_{\theta_{0}\in[\vartheta(\psi)\pm\frac{1}{\sqrt{n}}C]}\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}(\theta_{0})\right|\overset{p}{\longrightarrow}\inf_{\theta_{0}\in[\vartheta(\psi)\pm\frac{1}{\sqrt{n}}C]}\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t^{*}_{\Omega}(\theta_{0})\right|.$ (11)
Note that by the proof of Corollary A.3, $\left|t^{*}_{\Omega}\left(\vartheta(\psi)+\frac{1}{\sqrt{n}}\hat{\theta}^{*}_{\Sigma^{-1}}\right)\right|\leq\sqrt{J^{*}}$ for all $\Omega.$ Hence, for any $C\geq\left|\hat{\theta}^{*}_{\Sigma^{-1}}\right|,$ the right-hand side is bounded above by $\sqrt{J^{*},}$ regardless of $\kappa.$ Since $\sqrt{n}g_{n}(\psi)$ converges in distribution to a tight limit, so does $\hat{\theta}^{*}_{\Sigma^{-1}},$ so since the left hand side of (11) is decreasing in $C,$ and $\sqrt{J}-\sqrt{J^{*}}\overset{p}{\longrightarrow}0$ as shown in the proof of 5.2, we have that
$\limsup_{n\to\infty}\operatorname{P}\left\{\inf_{\theta_{0}}\sup_{\Omega\in\mathcal{W}_{k}}|t_{\Omega}(\theta_{0})|>\sqrt{J}+\varepsilon\right\}=0$ (12)

### v5

@@ 5.1. J-Statistics, not J-Tests
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}P\left(\left|\inf_{\theta_{0}\in\mathbb{R}}\sup_{\Omega\in\mathcal{W}_{\kappa}}|t_{\Omega}(\theta_{0})|-\sqrt{J}\right|>\varepsilon\right)=0.$
This result shows that, asymptotically, a researcher can find a weighting matrix that yields an absolute t-statistic of at least $\sqrt{J}$ for any possible null hypothesis. Hence, for instance, when $\sqrt{J}>1.96$ a nefarious researcher who is unconstrained in their choice of weighting matrix can reject any null hypothesis at the 5% level. Consequently, in order to limit the scope for “weight-hacking” to yield statistically significant results, it must be that either J-statistics are small or that other restrictions constrain the choice of weight matrix. [^28] Among other results, Honoreetal2020 characterize the derivative of $\sigma_{\Omega}^{2}$ with respect to elements of $\Omega.$ Their results may thus be useful for building intuition about which $\Omega$ satisfy $\hat{\sigma}_{\Omega}^{2}\leq(1+\tau^{2})\hat{\sigma}_{\Sigma^{-1}}^{2}$.
Absent restrictions on the weighting matrix, the J-statistic threshold necessary for a result to be “weight-hacking”-proof at the 5% level, $\sqrt{J}\leq 1.96$, does not scale with the degree of over-identification, and so can be much more stringent than conventional J-test critical values. In Table 1, for instance, four of ten specifications, and 3 of 6 papers, have $\sqrt{J}>1.96$, despite only one specification being rejected at the 5% level by the J-test.
@@
Under 5.1, for a critical value $c$, consider confidence sets centered around the GMM estimator $\hat{\theta}_{\Omega}$, $\hat{C}_{\Omega}(c)=[\hat{\theta}_{\Omega}\pm c\hat{\sigma}_{\Omega}]$. Then for all $\varepsilon>0,$
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}P\left(\left|\inf\left\{{c\geq 0:\bigcap_{\Omega\in\mathcal{W}_{\kappa}}C_{\Omega}(c)\neq\varnothing}\right\}-\sqrt{J}\right|>\varepsilon\right)=0.$
@@
$\lim_{\nu\to 0}\lim_{\kappa\to\infty}\limsup_{n\to\infty}P\left(\left|\sqrt{n}\cdot d_{H}\left(\bigcap_{\Omega\in\mathcal{W}_{\kappa}}C_{\Omega}(\sqrt{J}+\nu),\left\{\hat{\theta}_{\Sigma^{-1}}\right\}\right)\right|>\varepsilon\right)=0.$
Hence, to ensure at least one point of overlap between the confidence sets of researchers considering different weighting matrices, researchers would need to use critical values no smaller than $\sqrt{J}.$ This result bears an interesting resemblance to one by andrews2024pseudotrue. There, working in a Bayesian framework and studying linear-in-parameters moment condition models, the authors show that confidence intervals with width $\sqrt{J}$ have correct average coverage under a class of priors motivated by the GMM criterion.
Our asymptotic results in this section follow from exact, finite-sample results in a model, discussed in the Appendix, where the moments are linear in parameters and have known variance. armstrong2021sensitivity show that this linear model corresponds to the limit experiment for weakly identified GMM.
@@
Other recent work includes kitamura2013robustness, who establish a local-misspecification-robustness property for a generalized empirical likelihood estimator, and andrews2020informativeness who use a local misspecification framework to examine the relationship between structural estimates and other, non-moment statistics, with the goal of helping researchers clarify what “drives” their estimates. fessler2019shrinkage show how one can use the implications of locally misspecified economic models to improve unbiased but potentially noisy estimates, while armstrong2024adaptingmisspecification consider the problem of optimally combining biased and unbiased estimators under local misspecification. $\blacksquare$
@@ Appendix A Normal Limit Problem
armstrong2021sensitivity show that the problem of estimating $\theta$ under local misspecification reduces, asymptotically, to that of observing $(Y,\Gamma,\Sigma,h)$ where we wish to estimate $\theta=h^{\prime}\phi$ and
@@
Quantities associated with GMM have natural analogues in the limit problem (A), which we collect in Table 2. These quantities are linked in the sense that the asymptotic behavior of GMM quantities (left-side of Table 2) agrees with the behavior of the limit problem analogues under (A). In particular, GMM estimators for $\psi$ with (symmetric, positive definite) weight matrix $\Omega$ correspond to estimators of $\phi$ in the form
@@
$=(Y+\Gamma\hat{\phi}_{\Sigma^{-1}})\Sigma^{-1}(Y+\Gamma\hat{\phi}_{\Sigma^{-1}}).$
@@
Jacobian of moments, $E_{P}\left[{\frac{\partial{}}{\partial{\psi}}g(X_{i},\psi)}\right]$ $\Gamma$
@@
J-statistic $\min_{\psi\in\Psi}g_{n}(\psi)^{\prime}\operatorname{Var}_{P}(g(X_{i},\psi))^{-1}g_{n}(\psi)$ $\min_{u}(Y+\Gamma u)\Sigma^{-1}(Y+\Gamma u)$
@@
To clarify the limit problem (A), we briefly sketch the connection to the asymptotics of GMM. Suppose we are given some (nonrandom) $\tilde{\psi}$ with
$\phi\equiv\sqrt{n}(\tilde{\psi}-\psi)=O(1),$
that is, we are given a parameter value in a $\sqrt{n}$ neighborhood of the parameter of interest.
We next define a local parameter $\phi$, based on the deviation from $\tilde{\psi}$. Any estimator $\hat{\psi}$ of $\psi$ induces an estimator $\hat{\phi}$ of $\phi$ since
@@
$t_{\Omega}(\theta_{0})=\frac{h^{\prime}\hat{\phi}_{\Omega}-\theta_{0}}{\sqrt{\operatorname{Var}(h^{\prime}\hat{\phi}_{\Omega})}}.$
@@
Moreover, recall that $\hat{\theta}_{\Omega}=-h^{\prime}(\Gamma^{\prime}\Omega\Gamma)^{-1}\Gamma^{\prime}\Omega Y$, $\sigma_{\Omega}^{2}=\operatorname{Var}(\hat{\theta}_{\Omega})$. Then,
@@
    The J-statistic is the squared norm of $Z$: $J=Z^{\prime}\left({\Sigma_{Z}^{*}}\right)^{-1}Z$.
@@ A.1. Proofs for Normal Limit Problem
As a result, for $v_{\Omega(Z)}=c{\Sigma^{*}_{Z}}^{-1}Z$, we have that $\hat{\theta}_{\Omega(Z)}=\hat{\theta}_{\Sigma^{-1}}+cJ$ and
$\sigma_{\Omega(Z)}^{2}=\sigma_{\Sigma^{-1}}^{2}+c^{2}J.$
@@
$\hat{\theta}_{\Omega}-\hat{\theta}_{\Sigma^{-1}}=\sigma_{\Sigma^{-1}}\tau\sqrt{J}\quad\sigma_{\Omega}=\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}.$
@@
This is immediate from Corollary A.4. In particular, note that by Corollary A.4, $|t_{\Omega}(\hat{\theta}_{\Sigma^{-1}})|\leq\sqrt{J}$ for all $\Omega,$ while for any $\theta$ there exists $\Omega$ with $t_{\Omega}(\theta)\geq\sqrt{J}$. ∎
@@
To complete the proof of part (1) of the Proposition, the following is checked by Lemma A.6:
(8) $\left\{{\Omega\Gamma(\Gamma^{\prime}\Omega\Gamma)^{-1}h:\Omega}\right\}$ $=\left\{{q\in\mathbb{R}^{k}:\Gamma^{\prime}q=h}\right\}$
$=\left\{{q\in\mathbb{R}^{k}:q=q_{0}+v,\Gamma^{\prime}q_{0}=h,v\in N(\Gamma^{\prime})}\right\}.$
From this, it follows that the set of $v_{\Omega}$, as $\Omega$ ranges over all weighting matrices, is
$\left\{{-(M\Sigma M^{\prime})^{-1}M\Sigma q_{0}-(M\Sigma M^{\prime})^{-1}M\Sigma v:v\in N(\Gamma^{\prime})}\right\}.$
@@
($N(M)=R(\Gamma)$) $N(M\Sigma^{1/2})$ $=\left\{{v:M\Sigma^{1/2}v=0}\right\}=\left\{{v:\Sigma^{1/2}v\in R(\Gamma)}\right\}$
@@
$QY+Q\Gamma\hat{\phi}_{\Sigma^{-1}}$ $=\begin{bmatrix}\hat{\phi}_{\Sigma^{-1}}\\ Z\end{bmatrix}-\begin{bmatrix}\hat{\phi}_{\Sigma^{-1}}\\ 0\end{bmatrix}=\begin{bmatrix}0\\ Z\end{bmatrix}$
@@
Lemma A.6.
In the proof of A.5, (8) holds.
Proof.
$q=u+w$
for $P_{V}$ the projection matrix onto $V$. Note that by construction, any nonzero vector $t$ can be written as
$t=\Gamma t_{1}+cw+t_{2}\quad t_{1}\in\mathbb{R}^{p}\quad t_{2}\in V$
Set $a=\frac{1}{2h^{\prime}(\Gamma^{\prime}\Gamma)^{-2}h}>0$, and observe that $\Omega$ is positive definite:
(By Cauchy–Schwarz, $(t_{1}^{\prime}h)^{2}\leq(t_{1}^{\prime}(\Gamma^{\prime}\Gamma)^{2}t_{1})(h^{\prime}(\Gamma^{\prime}\Gamma)^{-2}h)$ ) $>a(t_{1}^{\prime}h)^{2}+ac^{2}\lVert w\rVert^{4}+2ac\lVert w\rVert^{2}t_{1}^{\prime}h+\lVert t_{2}\rVert^{2}$
$=a(t_{1}^{\prime}h+c\lVert w\rVert^{2})^{2}+\lVert t_{2}\rVert^{2}\geq 0.$
That the first inequality is strict when $t^{\prime}_{1}h\neq 0$ is straightforward. If $t^{\prime}_{1}h=0$, either $t_{1}\neq 0$ and Cauchy-Schwarz is strict or $t_{1}=0$, implying that $c$ or $t_{2}$ are not zero. In any case, the strict inequality in display is justified.
For the choice $a=\frac{1}{2h^{\prime}(\Gamma^{\prime}\Gamma)^{-2}h,}$ observe that
and thus
This proves (8). ∎
@@ Appendix B Proofs for Results in Main Text
(9) $\sqrt{n}(\hat{\theta}_{\Omega}-\vartheta(\psi))-\hat{\theta}^{*}_{\Omega}\overset{p}{\longrightarrow}0,~~\sqrt{n}\hat{\sigma}_{\Omega}\overset{p}{\longrightarrow}\sigma_{\Omega}=\left({h^{\prime}\left(\Gamma^{\prime}\Omega\Gamma\right)^{-1}\Gamma^{\prime}\Omega\Sigma\Omega\Gamma\left(\Gamma^{\prime}\Omega\Gamma\right)^{-1}h}\right)^{1/2},$
@@
$J=n\cdot g_{n}\left(\hat{\psi}_{\Sigma^{-1}}\right)^{\prime}\hat{\Sigma}^{-1}g_{n}\left(\hat{\psi}_{\Sigma^{-1}}\right)$
@@
$J^{*}=\min_{u}(\sqrt{n}g_{n}(\psi)+\Gamma u)^{\prime}\Sigma^{-1}(\sqrt{n}g_{n}(\psi)+\Gamma u),$
@@
$\sqrt{n}\cdot\left(\hat{\Theta}(\tau)-\vartheta(\psi)\right)=\left[\sqrt{n}(\hat{\theta}_{\Sigma^{-1}}-\vartheta(\psi))\pm\tau\sqrt{J}\sqrt{n}\hat{\sigma}_{\Sigma^{-1}}\right],$
@@
(10) $\text{plim}\left(d_{H}\left(\sqrt{n}\cdot\left(\hat{\Theta}\left(\kappa,\tau\right)-\vartheta(\psi)\right),\hat{\Theta}^{*}\left(\kappa,\tau\right)\right)\right)=0.$
@@
then $\lim_{\delta\to 0}d_{H}\left(\mathcal{W}_{\kappa,\tau},\mathcal{W}_{\kappa,\tau+\delta}\right)=0$ (where we now consider the Hausdorff distance defined based on e.g. Frobenius norm). Consequently, since (9) implies that for $\delta>0$
$P\left\{\mathcal{W}_{\kappa,\tau-\delta}\subseteq\mathcal{\hat{W}}_{\kappa,\tau}\subseteq\mathcal{W}_{\kappa,\tau+\delta}\right\}\to 1,$
@@
Our argument thus far implies that for sufficiently large $\kappa\geq 1$, and each $\nu>0,$
$\liminf_{n\to\infty}P\left(\sqrt{n}\cdot d_{H}\left(\hat{\Theta}\left(\kappa,\tau\right),\hat{\Theta}(\tau)\right)\leq d_{H}\left(\hat{\Theta}^{*}\left(\kappa,\tau\right),\hat{\Theta}^{*}\left(\tau\right)\right)+\nu\right)=1.$
To complete the proof, we wish to show that $d_{H}\left(\hat{\Theta}^{*}\left(\kappa,\tau\right),\hat{\Theta}^{*}\left(\tau\right)\right)\overset{p}{\longrightarrow}0$ as $\kappa\to\infty$. To this end, note that $\hat{\Theta}^{*}\left(\kappa,\tau\right)$ and $\hat{\Theta}^{*}\left(\tau\right)$ both depend on the data only through $\sqrt{n}g_{n}(\psi)$, and let
@@
$\liminf_{n\to\infty}P\left\{d_{H}\left(\hat{\Theta}^{*}\left(\kappa,\tau\right),\hat{\Theta}^{*}\left(\tau\right)\right)<\zeta\right\}\geq P\left\{\mathcal{N}\left(\eta,\Sigma\right)\in\mathcal{G}(\kappa,\zeta)\right\},$
@@
$\lim_{\kappa\to\infty}P\left\{N\left(\eta,\Sigma\right)\in\mathcal{G}(\kappa,\zeta)\right\}=1,$
@@
(11) $\inf_{\theta_{0}\in[\vartheta(\psi)\pm\frac{1}{\sqrt{n}}C]}\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}(\theta_{0})\right|\overset{p}{\longrightarrow}\inf_{\theta_{0}\in[\vartheta(\psi)\pm\frac{1}{\sqrt{n}}C]}\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t^{*}_{\Omega}(\theta_{0})\right|.$
Note that by the proof of Corollary A.3, $\left|t^{*}_{\Omega}\left(\vartheta(\psi)+\frac{1}{\sqrt{n}}\hat{\theta}^{*}_{\Sigma^{-1}}\right)\right|\leq\sqrt{J^{*}}$ for all $\Omega.$ Hence, for any $C\geq\left|\hat{\theta}^{*}_{\Sigma^{-1}}\right|,$ the right-hand side or (11) is bounded above by $\sqrt{J^{*},}$ regardless of $\kappa.$ Since $\sqrt{n}g_{n}(\psi)$ converges in distribution to a tight limit, so does $\hat{\theta}^{*}_{\Sigma^{-1}},$ so since the left hand side of (11) is decreasing in $C,$ and $\sqrt{J}-\sqrt{J^{*}}\overset{p}{\longrightarrow}0$ as shown in the proof of 5.2, we have that
(12) $\limsup_{n\to\infty}P\left\{\inf_{\theta_{0}}\sup_{\Omega\in\mathcal{W}_{k}}|t_{\Omega}(\theta_{0})|>\sqrt{J}+\varepsilon\right\}=0$